OpenStudy (gabylovesyou):

An original data set has the values 14, 2,11 , 6, and 2. What would need to be added to each value to change the mean to 15? 9 1 13 8 <------(correct answer??)) Please tell me if i'm right! :D

jimthompson5910 (jim_thompson5910):

sry i misread the problem

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (gabylovesyou):

ohhh ok yaay thank you!!!

jimthompson5910 (jim_thompson5910):

((14+x)+(2+x)+(11+x)+(6+x)+(2+x))/5 = 15 (35+5x)/5 = 15 35+5x = 15*5 35+5x = 75 5x = 75 - 35 5x = 40 x = 8 So this shows that you are correct

OpenStudy (gabylovesyou):

:D thank you

jimthompson5910 (jim_thompson5910):

you're welcome